# AI-Powered Inverse Design of Ku-Band SIW Resonant Structures by Iterative Residual Correction Network

### v1

@@ 2.3 Investigation of the Effect of Sweeping Design Parameters
To further understand the behavior of the proposed SIW multi-mode structure, a parametric analysis was performed by sweeping critical geometrical parameters and observing their impact on the filter’s electromagnetic response. Figures 4 and 5 illustrate the effect of varying the diameters $D_{1}$ and $D_{2}$, respectively, while keeping other parameters fixed.
Figure 4: Effect of varying via diameter $D_{1}$ on the S parameters of the multi-mode SIW structure.
Figure 5: Effect of varying via diameter $D_{2}$ on the S parameters of the multi-mode SIW structure.
The sweep results show that $D_{1}$ and $D_{2}$ significantly affect the center frequencies and bandwidths of the resonant modes. For example, increasing $D_{1}$ shifts the lower resonance, causing the resonances to move closer together. On the other hand, increasing $D_{2}$ results in the resonances moving farther apart.
Given that the coupling strength changes with variations in the distances $D_{1}$ and $D_{2}$, and consequently affects the $K$-inverters, it is natural and expected that these variations will influence the behavior of the S parameters.
While these sweeps offer valuable insight, they also highlight a key challenge in SIW filter design: the behavior of the structure is governed by multiple interdependent parameters. In this study, the full design is influenced not only by $D_{1}$ and $D_{2}$, but also by $R_{1}$, $R_{2}$, $R_{3}$, and $G$ — resulting in a six-dimensional parameter space. Each parameter can affect the resonant frequencies, bandwidth, and coupling strength.
As the number of design parameters grows, performing exhaustive electromagnetic simulations becomes computationally prohibitive. Moreover, traditional design methods relying on manual tuning or iterative sweeps are inefficient and often fail to capture subtle cross-parameter interactions.
To address this complexity, we propose using deep learning to model the forward and inverse relationship between SIW geometries and their electromagnetic responses. By generating a simulation-based dataset and training a neural network, we aim to:
  • Rapidly predict the design parameters for the arbitrary responses desired $S_{11}$ and $S_{21}$.
  • Perform inverse design, mapping desired frequency responses to optimal geometries.
  • Significantly reduce the computational cost and time compared to full-wave EM simulations.
@@ 3 Deep Learning Methodology
This study presents a deep learning-based framework for the inverse modeling of Ku-band SIW multimode resonators. Rather than relying on conventional single-pass or cascaded refinement strategies, the proposed approach utilizes an iterative residual correction mechanism to progressively enhance prediction accuracy through multiple feedback steps.
The overall deep learning strategy, as detailed in Section 3, is organized into four main components: dataset preparation, an initial feedforward inverse model (FIM), a hybrid inverse-forward residual refinement Network ((HiFR2-Net)), and a final iterative correction Network (IRC-Net).
@@ 3.1 Dataset Preparation
The dataset used in this study comprises 8,721 samples, each corresponding to a unique configuration of a SIW multimode resonator. Each configuration is defined by six geometric variables: $D_{1}$, $D_{2}$, $R_{1}$, $R_{2}$, $R_{3}$, and $G$, where $D_{1}$ and $D_{2}$ denote cavity dimensions, $R_{1}$, $R_{2}$, and $R_{3}$ represent via radius, and $G$ is the spacing between resonators. These parameters were generated using structured nested loops, constrained by physical and electromagnetic design rules to ensure valid realizations.
To enforce physically meaningful configurations, two constraints were applied during data generation:
$R_{3}\geq R_{2}\geq R_{1}\quad\text{and}\quad\left(\left(R_{1}+2R_{2}+D_{1}+D_{2}+2R_{3}\right)\cdot 2-0.4\right)/1.3<G.$
As mentioned earlier, the first constraint is used to realize the multi-mode resonator. Moreover, the second constraint ensures practical spacing between resonators and prevents overlapping via structures.
The sampling ranges were defined as follows:
@@
For each valid configuration, a full-wave electromagnetic simulation was performed using CST Studio Suite to extract the S-parameter response. Specifically, the magnitude responses of $S_{11}$ and $S_{21}$ were sampled across the Ku-band frequency range of 12–18 GHz at 1,001 frequency points, yielding a total of 2,002 input features per sample. The average simulation time per sample was approximately three minutes on a workstation equipped with an Intel Core i7-10750H CPU and 16 GB RAM, highlighting the computational cost of conventional methods.
The dataset preparation process was structured as follows:
@@
    Corresponding six-dimensional geometric parameters were collected and normalized, yielding the output matrix $\mathbf{Y}\in\mathbb{R}^{8721\times 6}$.
@@
    The dataset was randomly split into training and testing subsets with an 85%/15% ratio using a fixed seed (42) to ensure reproducibility.
This prepared dataset was used for training both the primary deep neural network model and the subsequent iterative residual correction models described in the following sections.
@@ 3.2 Feedforward Inverse Model
Table 2: Architecture of the Feedforward Inverse Network
@@ 3.3 Hybrid Inverse-Forward Residual Refinement Network
To enhance the prediction accuracy and consistency of inverse modeling, we introduce the Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net), schematically depicted in Figure 6. This cascaded learning framework integrates both inverse and forward learning stages, along with a residual refinement mechanism, to iteratively improve the fidelity of the design parameter estimation.
@@
  • Stage 1 – Feedforward Inverse Model (FIM): As previously detailed, the FIM serves as a baseline inverse predictor that generates an initial estimate $\mathbf{P}_{0}$ of the normalized geometric design parameters from the input S parameters.
  • Stage 2 – Feedforward Forward Model (FFM): The initial prediction $\mathbf{P}_{0}$ is fed into a forward model trained to reconstruct the original S-parameter vector $\hat{\mathbf{S}}$. This stage enforces forward consistency and enables indirect validation of the initial estimate through reconstruction error.
  • Stage 3 – Residual Refinement Model (RRM): The residual vector, computed as the difference between the true and reconstructed S parameters ($\mathbf{S}_{\text{true}}-\hat{\mathbf{S}}$), is used as input to a refinement model. This model predicts a correction term $\Delta\mathbf{P}$, which is added to the initial estimate $\mathbf{P}_{0}$ to improve alignment with the true design parameters.
@@
The integration of forward reconstruction and residual learning enables the system to better capture hidden dependencies and correct systematic estimation errors. In the next subsection, we further extend this idea by introducing an iterative correction mechanism to refine predictions through multiple passes.
Table 3: Architecture of the Feedforward Forward Network
@@
Input 6 (Dense) ReLU Estimated design parameters $\mathbf{P}_{0}$ from FIM
@@
Dropout 3 10% Regularization layer
Hidden 4 256 (Dense) ReLU Progressive feature expansion
Dropout 4 10% Regularization layer
@@
Output 2002 (Dense) ReLU Reconstructed S-parameter vector $\hat{\mathbf{S}}$
Figure 6: Schematic overview of the Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net).
@@ 3.4 Iterative Residual Correction Network
This modular architecture offers significant flexibility and interpretability. The FIM model captures the dominant inverse mapping, while the residual loop acts as a corrective mechanism to address subtle nonlinear discrepancies, enhancing generalization and robustness in the inverse modeling of complex electromagnetic structures.
while the iterative correction loop serves as an adaptive refinement mechanism.
@@
Table 4: Architecture of Residual Correction Network Used in Each Iteration
@@ 4 Results and Discussion
The section is organized into three main parts: (i) analysis of model learning behavior, focusing on the convergence and predictive capabilities of the FIM, the (HiFR2-Net), and the IRC-Net; (ii) statistical comparison of model performance using standard metrics; and (iii) practical validation through electromagnetic simulations and application-level implementations. This structure ensures a well-rounded assessment of both theoretical and practical aspects of the proposed methodology.
@@ 4.1.1 Learning Behavior of the FIM
Overall, the FIM demonstrates efficient and stable convergence with promising results in both training and validation phases. However, despite its effectiveness in providing an initial estimate, its performance can be further enhanced by employing an iterative refinement approach such as the Iterative Residual Correction Network (IRC-Net), which incrementally reduces prediction errors and improves model accuracy.
@@ 4.2 Statistical Evaluation
To provide a comprehensive understanding of model performance, we conducted a statistical analysis of prediction errors across the entire dataset, including training, validation, and test subsets. Two widely used error metrics, MSE and MAE, were employed to evaluate and compare the performance of the FIM, the HiFR2-Net, and the IRC-Net.
Figure 11 and Figure 12 illustrate the distribution of MAE and MSE values, respectively, on a logarithmic scale. These histograms provide insights not only into the average performance of each model but also into their consistency and robustness.
As shown in Figure 11, the MAE histogram reveals that the IRC-Net (green) achieves the lowest and most tightly clustered errors among all models. Its distribution peaks at a lower MAE value and exhibits minimal spread, indicating a high degree of consistency and fine-grained accuracy in its predictions. The HiFR2-Net (red) also demonstrates improved performance over the baseline FIM (blue), with a shifted peak toward lower errors and a narrower spread. In contrast, the FIM shows a broader distribution with a heavier tail on the right, reflecting its relatively higher prediction variability and larger occurrence of high-error instances.
Similarly, the MSE distribution in Figure 12 supports these observations. The IRC-Net maintains a compact error distribution with the majority of its predictions concentrated around the lowest MSE values. The HiFR2-Net again performs better than the FIM by achieving a more left-shifted peak and reduced variance. Notably, the logarithmic scale highlights the IRC-Net’s effectiveness in minimizing even small residual errors, a critical requirement in high-precision inverse design applications.
@@
Figure 11: Histogram of MAE on a logarithmic scale for FIM, HiFR2-Net, and IRC-Net. IRC-Net achieves the lowest and most consistent MAE values.
Figure 12: Histogram of Mean Squared Error MSE on a logarithmic scale for FIM, HiFR2-Net, and IRC-Net. The error distribution confirms the superior accuracy and reliability of IRC-Net.
4.3 Practical Validation through Simulation and Fabrication
To rigorously evaluate and validate the effectiveness of the proposed inverse design framework, we conducted both simulation-based and experimental assessments. Initially, a three-resonance SIW structure was designed and manufactured using conventional full-wave electromagnetic simulations to establish a baseline for training. This structure served as the foundation for the development and fine-tuning of the IRC-Net model. Subsequently, the trained IRC-Net was employed to predict the geometric parameters of a more complex four-resonance SIW structure, based solely on its desired frequency response.
The predicted designs for both the three- and four-resonance SIW components were then evaluated through full-wave electromagnetic simulations. As illustrated in Figure 13, the simulated S parameters demonstrate a strong agreement with the target responses, confirming the model’s ability to accurately capture and replicate the desired frequency behavior across multiple resonances within the Ku-band. Figure 14 displays the measurement setup used for validation, including the fabricated SIW resonator mounted between waveguide-to-coaxial adapters.
This superior performance highlights the practical value of the proposed IRC-Net framework for robust and precise inverse electromagnetic modeling. Not only does it provide substantial improvements over single-stage architectures such as the FIM, but also ensures physical realizability, as demonstrated through successful fabrication and testing of the designed structures. These findings validate the scalability and effectiveness of the proposed method and underscore its potential for advanced RF/microwave component design in real-world applications.

### v2

@@ 2.3 Investigation of the Effect of Sweeping Design Parameters
To further understand the behavior of the proposed SIW multi-mode structure, a parametric analysis was performed by sweeping critical geometrical parameters and observing their impact on the filter’s electromagnetic response. Figures 4(a) and 4(b) illustrate the effect of varying the diameters $D_{1}$ and $D_{2}$, respectively, while keeping other parameters fixed.
(a) Effect of varying via diameter $D_{1}$.
(b) Effect of varying via diameter $D_{2}$.
Figure 4: Effect of varying via diameters $D_{1}$ and $D_{2}$ on the S-parameters of the multi-mode SIW structure.
The sweep results indicate that both $D_{1}$ and $D_{2}$ have a significant impact on the center frequencies and bandwidths of the resonant modes. Specifically, increasing $D_{1}$ shifts the resonant frequencies to lower values and causes them to move closer together. Similarly, increasing $D_{2}$ also shifts the resonances to lower frequencies; however, it leads to an increased spacing between the resonant modes, resulting in the resonances moving farther apart.
Variations in distances $D_{1}$ and $D_{2}$ alter the coupling strength and, consequently, the characteristics of the $K$-inverters. As a result, these changes are expected to affect the behavior of the S-parameters.
Although the parameter sweeps provided useful insights into the roles of $D_{1}$ and $D_{2}$, they also underline a key limitation: as the number of design parameters increases, it becomes significantly more challenging to analyze the effects of each parameter individually. The interdependencies among parameters can lead to complex and sometimes unpredictable behavior, making it difficult to rely solely on parameter sweeps for accurate design optimization. In this study, six key parameters are considered in the design process: $D_{1}$, $D_{2}$, $R_{1}$, $R_{2}$, $R_{3}$, and $G$. Each of these parameters can significantly influence the resonant frequencies, bandwidth, and coupling strength.
To address the inherent complexity of this problem, deep learning is leveraged to model the forward and inverse relationships between SIW geometries and their corresponding electromagnetic characteristics. By generating a comprehensive simulation-based dataset and training a fully connected neural network, our approach aims to:
  • 1.
    Rapidly predict the design parameters that produce a broad range of customized responses in $S_{11}$ and $S_{21}$.
  • 2.
    Accurately perform inverse design, mapping desired frequency characteristics to optimal physical geometries.
  • 3.
    Significantly reduce the computational time and resource requirements compared to traditional full-wave electromagnetic simulations.
@@ 3 Deep Learning Methodology
This section outlines the deep learning framework developed for the inverse design of multimode SIW filters. The methodology is structured into four key stages, each building upon the previous to progressively enhance prediction accuracy and robustness. The process begins with Dataset Preparation, which details the generation and preprocessing of input-output pairs—comprising S-parameters and design parameters—for neural network training. Next, the Feedforward Inverse Model (FIM)is introduced as the foundational network that estimates the filter design based on measured S-parameters. Building on this foundation, the Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net) is presented, incorporating both inverse and forward mappings to refine predictions through residual correction. Finally, the Iterative Residual Correction Network (IRC-Net) is described, employing successive correction loops to iteratively adjust the design estimate, thereby aligning it more closely with the true design parameters.
This progressive approach demonstrates how deep learning architectures can be adapted and enhanced to solve complex inverse design problems with increasing levels of precision and robustness.
@@ 3.1 Dataset Preparation
The dataset used in this study comprises 8,721 samples, each corresponding to a unique configuration of an SIW multimode resonator. Each configuration is defined by six geometric parameters: $D_{1}$, $D_{2}$, $R_{1}$, $R_{2}$, $R_{3}$, and $G$. These parameters were generated through structured nested loops, constrained by physical and electromagnetic design rules to ensure valid configurations.
Two constraints were applied during data generation to guarantee physically meaningful and practical designs:
$R_{3}\geq R_{2}\geq R_{1},\quad\text{and}\quad\frac{\left(R_{1}+2R_{2}+D_{1}+D_{2}+2R_{3}\right)\times 2-0.4}{1.3}<G.$
The first constraint ensures proper multi-mode resonator construction, while the second enforces adequate spacing between resonators to prevent overlapping via structures.
Sampling ranges for the parameters were defined as follows:
  • 1.
@@
For each valid configuration, full-wave electromagnetic simulations were performed using CST Studio Suite to extract the S-parameter responses. The magnitude responses of $S_{11}$ and $S_{21}$ were sampled across the Ku-band frequency range of 9–20 GHz at 1,001 frequency points, resulting in a total of 2,002 input features per sample.
The entire dataset was generated over approximately 92 hours on a workstation equipped with a 13th Gen Intel(R) Core(TM) i5-13600K CPU running at 3.5 GHz and 128 GB of RAM, highlighting the significant computational cost of conventional simulation methods.
Dataset preparation involved the following steps:
@@
    Corresponding six-dimensional geometric parameters were collected and normalized, forming the output matrix $\mathbf{Y}\in\mathbb{R}^{8721\times 6}$.
@@
    The dataset was split into 80% for training and 20% for testing, using a fixed random seed (42) to guarantee reproducibility.
This prepared dataset was utilized for training both the primary deep neural network and the subsequent iterative residual correction models detailed in the following sections. Dataset preprocessing and model training were conducted on the Google Colab platform, where data preparation scripts leveraged Colab’s computational resources for efficient processing, and GPU acceleration significantly expedited model convergence on the large-scale dataset.
@@ 3.2 Feedforward Inverse Model
Table 3
Architecture of the Feedforward Inverse Network.
@@ 3.3 Hybrid Inverse-Forward Residual Refinement Network
To enhance the prediction accuracy and consistency of inverse modeling, the Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net) is introduced, as schematically depicted in Figure 5. This cascaded learning framework integrates both inverse and forward learning stages, along with a residual refinement mechanism, to iteratively improve the fidelity of the design parameter estimation.
@@
  • 1.
    Stage 1 – Feedforward Inverse Model: As previously detailed, the FIM serves as a baseline inverse predictor that generates an initial estimate $\mathbf{P}_{0}$ of the normalized geometric design parameters from the input S parameters.
  • 2.
    Stage 2 – Feedforward Forward Model: The initial prediction $\mathbf{P}_{0}$ is fed into a forward model trained to reconstruct the original S-parameter vector $\hat{\mathbf{S}}$. This stage enforces forward consistency and enables indirect validation of the initial estimate through reconstruction error.
  • 3.
    Stage 3 – Residual Refinement Model: The residual vector, computed as the difference between the true and reconstructed S parameters ($\mathbf{S}_{\text{true}}-\hat{\mathbf{S}}$), is used as input to a refinement model. This model predicts a correction term $\Delta\mathbf{P}$, which is added to the initial estimate $\mathbf{P}_{0}$ to improve alignment with the true design parameters.
@@
The integration of forward reconstruction and residual learning enables the system to better capture hidden dependencies and correct systematic estimation errors. The next subsection extends this concept by introducing an iterative correction mechanism that refines predictions through multiple passes.
Table 4
Architecture of the Feedforward Forward Network.
@@
Input 6 (Dense) ReLU Estimated parameters $\mathbf{P}_{0}$ from FIM
@@
Dropout 3 10% Regularization
Hidden 4 256 (Dense) ReLU Feature expansion
Dropout 4 10% Regularization
@@
Output 2002 (Dense) ReLU Reconstructed output vector $\hat{\mathbf{S}}$
Figure 5: Schematic overview of the Hybrid Inverse-Forward Residual Refinement Network.
@@ 3.4 Iterative Residual Correction Network
This modular architecture provides both flexibility and interpretability. The FIM captures the dominant inverse mapping, while the residual correction loop functions as an adaptive refinement mechanism to address subtle nonlinear discrepancies—thereby improving generalization and robustness in the inverse modeling of complex electromagnetic structures.
@@
Table 5
Architecture of Residual Correction Network used in each iteration.
@@ 4 Results and Discussion
The section is organized into three main parts: (i) analysis of model learning behavior, focusing on the convergence and predictive capabilities of the FIM, HiFR2-Net, and the IRC-Net; (ii) statistical comparison of model performance using standard metrics; and (iii) practical validation through electromagnetic simulations and application-level implementations. This structure ensures a well-rounded assessment of both theoretical and practical aspects of the proposed methodology.
In order to support the practical viability of the proposed models, we also report the computational cost and training time of each approach. All models were trained on Google Colab using an NVIDIA Tesla T4 GPU, which provides 16 GB of GDDR6 memory and is optimized for both training and inference workloads. The baseline models (FIM and HiFR2-Net) were trained with a batch size of 128 using early stopping, while the iterative correction stage in IRC-Net employed a batch size of 32 for 100 epochs. The total training durations were approximately 161 seconds for FIM, 477 seconds for HiFR2-Net, and 867 seconds for IRC-Net (including 160 seconds for the initial FIM and 707 seconds for 5 iterations of residual correction). The prediction time per sample for all models is less than one millisecond, confirming the suitability of the proposed framework for real-time applications.
@@ 4.1.1 Learning Behavior of the FIM
Overall, the FIM demonstrates efficient and stable convergence with promising results in both training and validation phases. However, despite its effectiveness in providing an initial estimate, its performance can be further enhanced by employing an iterative refinement approach such as the IRC-Net, which incrementally reduces prediction errors and improves model accuracy.
@@ 4.2 Statistical Evaluation
To provide a comprehensive understanding of model performance, a statistical analysis of prediction errors was conducted across the training, validation, and test subsets. Two widely used error metrics, MSE and MAE, were employed to evaluate and compare the performance of the FIM, the HiFR2-Net, and the IRC-Net.
Figure 10(a) and Figure 10(b) illustrate the distribution of MAE and MSE values, respectively, on a logarithmic scale. These histograms provide insights not only into the average performance of each model but also into their consistency and robustness.
As shown in Figure 10(a), the MAE histogram reveals that the IRC-Net (green) achieves the lowest and most tightly clustered errors among all models. Its distribution peaks at a lower MAE value and exhibits minimal spread, indicating a high degree of consistency and fine-grained accuracy in its predictions. The HiFR2-Net (red) also demonstrates improved performance over the baseline FIM (blue), with a shifted peak toward lower errors and a narrower spread. In contrast, the FIM shows a broader distribution with a heavier tail on the right, reflecting its relatively higher prediction variability and larger occurrence of high-error instances.
Similarly, the MSE distribution in Figure 10(b) supports these observations. The IRC-Net maintains a compact error distribution with the majority of its predictions concentrated around the lowest MSE values. The HiFR2-Net again performs better than the FIM by achieving a more left-shifted peak and reduced variance. Notably, the logarithmic scale highlights the IRC-Net’s effectiveness in minimizing even small residual errors, a critical requirement in high-precision inverse design applications.
@@
(a) Histogram of MAE (log scale) for FIM, HiFR2-Net, and IRC-Net.
(b) Histogram of MSE (log scale) for FIM, HiFR2-Net, and IRC-Net.
Figure 10: Log-scale histograms of MAE and MSE for three network architectures.
4.3 Parametric Evaluation and Experimental Validation
To comprehensively evaluate the performance of the proposed IRC-Net framework, its capability to accurately predict the design parameters of multimode SIW resonators is first assessed. The model’s predictions are compared against the ground truth for the structure illustrated in Figure 3, which was randomly selected from the test dataset. The predicted design parameters from the three models, FIM, HiFR2-Net, and IRC-Net are compared with the true target values in Table 6, where the IRC-Net demonstrates the closest match to the target parameters, highlighting its superior accuracy in inverse modeling.
Table 6
Comparison between model predictions and ground truth for various parameters.
Parameters FIM-Net HiFR2-Net IRC-Net Ground Truth
D1 5.5871 5.8126 5.5745 5.5000
D2 7.8192 7.9293 7.9494 8.0000
R1 0.2118 0.2136 0.2174 0.2000
R2 0.3846 0.3723 0.4084 0.4000
R3 0.8210 0.8065 0.8047 0.8000
G 25.7656 26.1080 26.0227 26.0000
Table 6 clearly demonstrates that the IRC-Net model provides significantly more accurate predictions of the design parameters compared to the FIM and HiFR2-Net models. To further validate these findings, simulations were performed using the predicted design parameters. Only the resulting $S_{21}$ responses are presented for streamlined comparison, as shown in Figure 11. The simulations provide an empirical validation by comparing electromagnetic performance against the ground truth, and the IRC-Net yields an $S_{21}$ response closely matching the ground truth across the entire frequency range, outperforming both the FIM and HiFR2-Net models.
Figure 11: Simulated $S_{21}$ parameter responses using predicted design parameters from different models compared with the ground truth.
To evaluate the practical effectiveness of the proposed IRC-Net framework, multi-mode SIW resonator structures were selected for fabrication and experimental validation. The first design, which exhibits a four-resonance response, was developed using conventional full-wave electromagnetic simulation techniques. In contrast, the second structure, which features three in-band resonances, was designed entirely by the trained IRC-Net model based on the desired frequency response. Both structures were fabricated using standard PCB manufacturing processes on the RT/Duroid 5880 substrate with a thickness of 20 mils. The S-parameters of the fabricated devices were measured using an HP8510 vector network analyzer, as illustrated in Figure 12.
(a) Fabricated SIW structure exhibiting three in-band resonances.
(b) Fabricated SIW structure exhibiting four in-band resonances.
Figure 12: Measurement setup for the fabricated SIW structures using an HP8510 VNA.
The input signal was generated by an external frequency synthesizer, and the input power level was set to $30\,\mathrm{dBm}$ to ensure that the input mixers operate within their linear range. The frequency sweep was performed from 9 GHz to 20 GHz.
Figure 13 presents a comparison between the measured, simulated, and predicted S-parameters. As observed, there is an excellent agreement between the fabricated and predicted results, confirming the effectiveness of the proposed model. This strong correlation indicates that the IRC-Net framework can be reliably extended to more complex passive microwave resonant structures involving higher design freedom and a larger number of geometric parameters. In summary, the fabrication and measurement results of both the three- and four-resonance SIW prototypes validate the high accuracy and practical applicability of the proposed framework, demonstrating its capability to guide the design of complex multi-mode resonators with confidence.